# Accelerating Quantum Reinforcement Learning with a Quantum Natural Policy Gradient Based Approach

### v1

@@ Appendix E Proof of Lemma 6
$\leq 3\hat{\delta}_{F}^{2}\mathbb{E}\left[\|\omega_{h}-\omega_{k}^{*}\|^{2}\right]+\frac{3\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}\hat{\delta}_{F}^{2}+3\hat{\delta}_{g}^{2}$
@@
$\overset{(a)}{\leq}3\hat{\delta}_{F}^{2}\bigg{\{}\mathbb{E}\left[\|\omega_{0}-\omega_{k}^{*}\|^{2}\right]+\alpha R_{0}+R_{1}\bigg{\}}+\frac{3\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}\hat{\delta}_{F}^{2}+3\hat{\delta}_{g}^{2}$
@@
$-\langle\mathbb{E}[\omega_{h}]$ $-\omega_{k}^{*},\mathbb{E}\left[\mathbb{E}[\nabla_{\omega}\tilde{L}_{h}]-F_{\rho}(\theta_{k})(\mathbb{E}[\omega_{h}]-\omega_{k}^{*})\right]\rangle$
$\leq\frac{\mu_{F}}{4}\|\mathbb{E}[\omega_{h}]-\omega_{k}^{*}\|^{2}+\frac{1}{\mu_{F}}\left\|\mathbb{E}[\nabla_{\omega}\tilde{L}_{h}]-F_{\rho}(\theta_{k})(\mathbb{E}[\omega_{h}]-\omega_{k}^{*})\right\|^{2}$
$\leq\frac{\mu_{F}}{4}\|\mathbb{E}[\omega_{h}]-\omega_{k}^{*}\|^{2}+\dfrac{3}{\mu_{F}}\left[\hat{\delta}_{F}^{2}\bigg{\{}\mathbb{E}\left[\|\omega_{0}-\omega_{k}^{*}\|^{2}\right]+\alpha R_{0}+R_{1}\bigg{\}}+\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}\hat{\delta}_{F}^{2}+\hat{\delta}_{g}^{2}\right]$
@@
$\|\mathbb{E}[\omega_{h+1}]-\omega_{k}^{*}\|^{2}$ $\leq\left(1-\dfrac{3\alpha\mu_{F}}{2}+2G^{4}\alpha^{2}\right)\|\mathbb{E}[\omega_{h}]-\omega_{k}^{*}\|^{2}$
$\hskip 28.45274pt+6\alpha\left(\alpha+\dfrac{1}{\mu_{F}}\right)\left[\hat{\delta}_{F}^{2}\bigg{\{}\mathbb{E}\left[\|\omega_{0}-\omega_{k}^{*}\|^{2}\right]+\alpha R_{0}+R_{1}\bigg{\}}+\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}\hat{\delta}_{F}^{2}+\hat{\delta}_{g}^{2}\right]$
@@
$\|\mathbb{E}[\omega_{h+1}]-\omega_{k}^{*}\|^{2}$ $\leq\left(1-{\alpha\mu_{F}}\right)\|\mathbb{E}[\omega_{h}]-\omega_{k}^{*}\|^{2}$
$\hskip 28.45274pt+6\alpha\left(\alpha+\dfrac{1}{\mu_{F}}\right)\left[\hat{\delta}_{F}^{2}\bigg{\{}\mathbb{E}\left[\|\omega_{0}-\omega_{k}^{*}\|^{2}\right]+\alpha R_{0}+R_{1}\bigg{\}}+\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}\hat{\delta}_{F}^{2}+\hat{\delta}_{g}^{2}\right]$
@@
$\|\mathbb{E}[\omega_{H}]-\omega_{k}^{*}\|^{2}\leq\left(1-{\alpha\mu_{F}}\right)^{H}\|\mathbb{E}[\omega_{0}]-\omega_{k}^{*}\|^{2}$
$\hskip 42.67912pt+\sum_{h=0}^{H-1}6\left(1-{\alpha\mu_{F}}\right)^{h}\alpha\left(\alpha+\dfrac{1}{\mu_{F}}\right)\left[\hat{\delta}_{F}^{2}\bigg{\{}\mathbb{E}\left[\|\omega_{0}-\omega_{k}^{*}\|^{2}\right]+\alpha R_{0}+R_{1}\bigg{\}}+\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}\hat{\delta}_{F}^{2}+\hat{\delta}_{g}^{2}\right]$
$\leq\exp\left(-{H\alpha\mu_{F}}\right)\|\mathbb{E}[\omega_{0}]-\omega_{k}^{*}\|^{2}+\dfrac{6}{\mu_{F}}\left(\alpha+\dfrac{1}{\mu_{F}}\right)\left[\hat{\delta}_{F}^{2}\bigg{\{}\mathbb{E}\left[\|\omega_{0}-\omega_{k}^{*}\|^{2}\right]+\alpha R_{0}+R_{1}\bigg{\}}+\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}\hat{\delta}_{F}^{2}+\hat{\delta}_{g}^{2}\right]$
@@ Appendix F Proof of Theorem 3
$\begin{split}J_{\rho}^{*}-\dfrac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[J_{\rho}(\theta_{k})]&\leq\sqrt{\epsilon_{\mathrm{bias}}}+G\exp\left(-{\frac{H\alpha\mu_{F}}{2}}\right)\|\mathbb{E}[\omega_{0}]-\omega^{*}\|^{2}+C_{2}\\ &+\dfrac{B}{4L}\left(\dfrac{\mu_{F}^{2}}{G^{2}}+G^{2}\right)\exp\left(-{H\alpha\mu_{F}}\right)\mathbb{E}\|\omega_{0}-\omega^{*}\|^{2}+C_{3}+C_{4}\end{split}$ (64)
with $C_{2}=G\sqrt{C_{1}}$, $C_{3}=\dfrac{B}{4L}\left(\dfrac{\mu_{F}^{2}}{G^{2}}+G^{2}\right)C_{0}$ and $C_{4}=\dfrac{G^{2}}{\mu_{F}^{2}K}\left(\dfrac{B}{1-\gamma}+4L\mathbb{E}_{s\sim d_{\rho}^{\pi^{*}}}[KL(\pi^{*}(\cdot|s)\|\pi_{\theta_{0}}(\cdot|s))]\right)$. Where $C_{0}$ and $C_{1}$ are defined in Lemma 6. By replacing $C_{4}=\operatorname*{\mathcal{O}}{(\frac{1}{K})}$ we can express (64) as follows:
$J_{\rho}^{*}-\dfrac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[J_{\rho}(\theta_{k})]\leq\sqrt{\epsilon_{\mathrm{bias}}}+\operatorname*{\mathcal{O}}(e^{-H})+{\operatorname*{\mathcal{O}}}\left(\gamma^{2N}+\frac{\hat{\sigma}_{F}^{2}}{(1-\gamma)^{4}}+\hat{\sigma}_{g}^{2}\right)+\operatorname*{\mathcal{O}}{(\frac{1}{K})}$ (65)

### v3

@@ Appendix E Proof of Lemma 6
$\leq 3\hat{\delta}_{F}^{2}\operatorname{\mathbb{E}}\left[\|\omega_{h}-\omega_{k}^{*}\|^{2}\right]+\frac{3\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}\hat{\delta}_{F}^{2}+3\hat{\delta}_{g}^{2}$
@@
$\overset{(a)}{\leq}3\hat{\delta}_{F}^{2}\bigg{\{}\operatorname{\mathbb{E}}\left[\|\omega_{0}-\omega_{k}^{*}\|^{2}\right]+\alpha R_{0}+R_{1}\bigg{\}}+\frac{3\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}\hat{\delta}_{F}^{2}+3\hat{\delta}_{g}^{2}$
@@
$-\langle\operatorname{\mathbb{E}}[\omega_{h}]$ $-\omega_{k}^{*},\operatorname{\mathbb{E}}\left[\operatorname{\mathbb{E}}[\nabla_{\omega}\tilde{L}_{h}]-F_{\rho}(\theta_{k})(\operatorname{\mathbb{E}}[\omega_{h}]-\omega_{k}^{*})\right]\rangle$
$\leq\frac{\mu_{F}}{4}\|\operatorname{\mathbb{E}}[\omega_{h}]-\omega_{k}^{*}\|^{2}+\frac{1}{\mu_{F}}\left\|\operatorname{\mathbb{E}}[\nabla_{\omega}\tilde{L}_{h}]-F_{\rho}(\theta_{k})(\operatorname{\mathbb{E}}[\omega_{h}]-\omega_{k}^{*})\right\|^{2}$
$\leq\frac{\mu_{F}}{4}\|\operatorname{\mathbb{E}}[\omega_{h}]-\omega_{k}^{*}\|^{2}+\dfrac{3}{\mu_{F}}\left[\hat{\delta}_{F}^{2}\bigg{\{}\operatorname{\mathbb{E}}\left[\|\omega_{0}-\omega_{k}^{*}\|^{2}\right]+\alpha R_{0}+R_{1}\bigg{\}}+\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}\hat{\delta}_{F}^{2}+\hat{\delta}_{g}^{2}\right]$
@@
$\|\operatorname{\mathbb{E}}[\omega_{h+1}]-\omega_{k}^{*}\|^{2}$ $\leq\left(1-\dfrac{3\alpha\mu_{F}}{2}+2G^{4}\alpha^{2}\right)\|\operatorname{\mathbb{E}}[\omega_{h}]-\omega_{k}^{*}\|^{2}$
$\hskip 28.45274pt+6\alpha\left(\alpha+\dfrac{1}{\mu_{F}}\right)\left[\hat{\delta}_{F}^{2}\bigg{\{}\operatorname{\mathbb{E}}\left[\|\omega_{0}-\omega_{k}^{*}\|^{2}\right]+\alpha R_{0}+R_{1}\bigg{\}}+\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}\hat{\delta}_{F}^{2}+\hat{\delta}_{g}^{2}\right]$
@@
$\|\operatorname{\mathbb{E}}[\omega_{h+1}]-\omega_{k}^{*}\|^{2}$ $\leq\left(1-{\alpha\mu_{F}}\right)\|\operatorname{\mathbb{E}}[\omega_{h}]-\omega_{k}^{*}\|^{2}$
$\hskip 28.45274pt+6\alpha\left(\alpha+\dfrac{1}{\mu_{F}}\right)\left[\hat{\delta}_{F}^{2}\bigg{\{}\operatorname{\mathbb{E}}\left[\|\omega_{0}-\omega_{k}^{*}\|^{2}\right]+\alpha R_{0}+R_{1}\bigg{\}}+\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}\hat{\delta}_{F}^{2}+\hat{\delta}_{g}^{2}\right]$
@@
$\|\operatorname{\mathbb{E}}[\omega_{H}]-\omega_{k}^{*}\|^{2}\leq\left(1-{\alpha\mu_{F}}\right)^{H}\|\operatorname{\mathbb{E}}[\omega_{0}]-\omega_{k}^{*}\|^{2}$
$\hskip 42.67912pt+\sum_{h=0}^{H-1}6\left(1-{\alpha\mu_{F}}\right)^{h}\alpha\left(\alpha+\dfrac{1}{\mu_{F}}\right)\left[\hat{\delta}_{F}^{2}\bigg{\{}\operatorname{\mathbb{E}}\left[\|\omega_{0}-\omega_{k}^{*}\|^{2}\right]+\alpha R_{0}+R_{1}\bigg{\}}+\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}\hat{\delta}_{F}^{2}+\hat{\delta}_{g}^{2}\right]$
$\leq\exp\left(-{H\alpha\mu_{F}}\right)\|\operatorname{\mathbb{E}}[\omega_{0}]-\omega_{k}^{*}\|^{2}+\dfrac{6}{\mu_{F}}\left(\alpha+\dfrac{1}{\mu_{F}}\right)\left[\hat{\delta}_{F}^{2}\bigg{\{}\operatorname{\mathbb{E}}\left[\|\omega_{0}-\omega_{k}^{*}\|^{2}\right]+\alpha R_{0}+R_{1}\bigg{\}}+\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}\hat{\delta}_{F}^{2}+\hat{\delta}_{g}^{2}\right]$
@@ Appendix F Proof of Theorem 3
$\begin{split}J_{\rho}^{*}-\dfrac{1}{K}\sum_{k=0}^{K-1}\operatorname{\mathbb{E}}[J_{\rho}(\theta_{k})]&\leq\sqrt{\epsilon_{\mathrm{bias}}}+G\exp\left(-{\frac{H\alpha\mu_{F}}{2}}\right)\|\operatorname{\mathbb{E}}[\omega_{0}]-\omega^{*}\|^{2}+C_{2}\\ &+\dfrac{B}{4L}\left(\dfrac{\mu_{F}^{2}}{G^{2}}+G^{2}\right)\exp\left(-{H\alpha\mu_{F}}\right)\operatorname{\mathbb{E}}\|\omega_{0}-\omega^{*}\|^{2}+C_{3}+C_{4}\end{split}$ (64)
with $C_{2}=G\sqrt{C_{1}}$, $C_{3}=\dfrac{B}{4L}\left(\dfrac{\mu_{F}^{2}}{G^{2}}+G^{2}\right)C_{0}$ and $C_{4}=\dfrac{G^{2}}{\mu_{F}^{2}K}\left(\dfrac{B}{1-\gamma}+4L\operatorname{\mathbb{E}}_{s\sim d_{\rho}^{\pi^{*}}}[KL(\pi^{*}(\cdot|s)\|\pi_{\theta_{0}}(\cdot|s))]\right)$. Where $C_{0}$ and $C_{1}$ are defined in Lemma 6. By replacing $C_{4}=\operatorname*{\mathcal{O}}{(\frac{1}{K})}$ we can express (64) as follows:
$J_{\rho}^{*}-\dfrac{1}{K}\sum_{k=0}^{K-1}\operatorname{\mathbb{E}}[J_{\rho}(\theta_{k})]\leq\sqrt{\epsilon_{\mathrm{bias}}}+\operatorname*{\mathcal{O}}(e^{-H})+{\operatorname*{\mathcal{O}}}\left(\gamma^{2N}+\frac{\hat{\sigma}_{F}^{2}}{(1-\gamma)^{4}}+\hat{\sigma}_{g}^{2}\right)+\operatorname*{\mathcal{O}}{(\frac{1}{K})}$ (65)